Circular No. 7064

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET P/1939 TN = 1998 WG_{22} (VÄISÄLÄ-OTERMA) Precise positions of 1939 TN, an apparently asteroidal object discov-

ered by Y. Väisälä at Turku on 1939 Oct. 7 and observed by him and L. Oterma on three more nights over a 35-day arc, were published in 1979 in Turku Obs. Report R10 and on MPC 4811. The orbit computation by Oterma suggested to her that the object was a comet, and in a communication to the Minor Planet Center in 1981 she remarked that a careful reexamination of the plates suggested that the object was perhaps somewhat diffuse. The introduction to the 1982 edition of the Catalogue of Cometary Orbits quotes her conclusion that the object was probably a comet, but the object was not actually listed as one. On MPEC 1998-X19, the object is identified by S. Nakano, Sumoto, with 1998 WG_{22}, an apparently asteroidal object observed by the LINEAR program on Nov. 18 and 21:

1998 UT  α_{2000}  δ_{2000}  m_{2}
Nov. 18.25966  3^{h}40^{m}20.^{s}80  +19^{o}53′48′′.8  19.0

Fortuitously, this object was located only 4′ from the result of integrating forward from 1939 the orbital elements by the undersigned on MPC 6815. Further LINEAR observations were made on Nov. 24. On observing the object with the 1.8-m reflector at the Dominion Astrophysical Observatory on Dec. 6.3 UT, D. Balam noted (in FWHM 3′′.1 seeing) an 8′′ coma and a tail extending 18′′ in p.a. 260^{o}. On Dec. 7.1 W. Offutt, Cloudcroft, observing between clouds, also remarked on a tail (or antitail) extending 17′′ in p.a. 257^{o}, but no coma was noted. M. Tichý, observing in poor conditions at Kleť on Dec. 7.9, remarked on a possible 6′′ coma.

The object’s cometary nature seems now reasonably assured, and the following are Nakano’s elements on MPEC 1998-X19 in “cometary” style:

Epoch = 1999 Jan. 22.0 TT

T = 1998 Sept.27.2636 TT  ω = 165.^{o}3358
e = 0.246318  Ω = 242.4572 }2000.0
q = 3.391581 AU  i = 2.3399

a = 4.500018 AU n^{o} = 0.1032483 P = 9.546 years

1998/99 α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Dec. 3 3^{h}30.^{m}68 +19^{o}11.^{′}6 2.446 3.406 164.^{o}4 4.^{o}5 16.8
13 3 25.25 +18 45.9 2.504 3.410 152.9  7.5 16.8
23 3 21.35 +18 25.6 2.588 3.415 141.8 10.3 16.9
Jan. 2 3 19.34 +18 12.5 2.694 3.421 131.1 12.5 17.0
1998 December 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden